### v1

# From Chat Logs to Collective Insights: Aggregative Question Answering

Abstract
Conversational agents powered by large language models (LLMs) are rapidly becoming integral to our daily interactions, generating unprecedented amounts of conversational data. Such datasets offer a powerful lens into societal interests, trending topics, and collective concerns. Yet, existing approaches typically treat these interactions as independent and miss critical insights that could emerge from aggregating and reasoning across large-scale conversation logs. In this paper, we introduce Aggregative Question Answering, a novel task requiring models to reason explicitly over thousands of user-chatbot interactions to answer aggregative queries, such as identifying emerging concerns among specific demographics. To enable research in this direction, we construct a benchmark, WildChat-AQA, comprising 6,027 aggregative questions derived from 182,330 real-world chatbot conversations. Experiments show that existing methods either struggle to reason effectively or incur prohibitive computational costs, underscoring the need for new approaches capable of extracting collective insights from large-scale conversational data.
[…]
From Chat Logs to Collective Insights:
Aggregative Question Answering
Wentao Zhang University of Waterloo w564zhan@uwaterloo.ca Woojeong Kim Cornell University wk247@cornell.edu Yuntian Deng University of Waterloo yuntian@uwaterloo.ca
1 Introduction
Rapid adoption of conversation agents powered by large language models (LLMs) is transforming human-computer interactions, integrating deeply into society, and generating unprecedented volumes of conversational data (Backlinko Team, 2025; Vynck, 2023). Platforms using LLM-based chatbots now routinely handle millions of interactions daily, producing rich datasets that capture real-time dialogues reflecting genuine user interests, emerging societal trends, and collective concerns (Zhao et al., 2024b; Zheng et al., 2024). Such conversational data offer immense potential for deriving insights at scale, revealing patterns in societal dynamics, shifts in public sentiment, and demographic-specific concerns.
Despite the inherent richness of these conversational datasets, current research typically treats interactions as isolated, independent data points, primarily using them to finetune LLMs for generating improved individual responses (The Vicuna Team, 2023; Lambert et al., 2025; Zhang et al., 2025). This independent and identically distributed (i.i.d.) assumption overlooks important temporal patterns and thematic connections that naturally arise from large-scale, real-world user-chatbot conversations. Conversations do not occur in isolation, but rather within specific temporal, geographical, and device-related contexts (Tamkin et al., 2024). These contextual features carry significant potential for deriving collective insights, such as understanding regional differences in user concerns or identifying temporal shifts in societal attitudes—insights which are lost under the simplifying i.i.d. assumption.
To address this gap, we introduce a new task, Aggregative Question Answering, which requires reasoning across large-scale collections of user-chatbot interactions to extract aggregative insights. Unlike traditional summarization, which condenses information from one or a few documents into static summaries, Aggregative Question Answering generates dynamic answers that explicitly depend on the specific aggregative query posed. The task requires holistic reasoning over thousands of conversations to answer questions such as identifying trending topics within specific timeframes (What topics trended last week?”), emerging concerns among particular demographics (What topics are Californians concerned about before an election?”), or tracking changes in societal sentiment (“How have users’ attitudes toward artificial intelligence evolved this month?”). The core challenge thus lies not in summarizing individual conversations, but rather in global-scale reasoning conditioned on the query. Figure 1 highlights the high-level distinctions between traditional summarization, aggregative question answering, and querying predefined databases.
To facilitate research into Aggregative Question Answering, we introduce a benchmark WildChat-AQA, constructed from the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat captures not only conversation transcripts but also metadata such as temporal, geographical, and user-specific information. WildChat-AQA formulates aggregational queries about both explicit and implicit attributes of conversations—including topics, keywords, geographical locations, and time—in a multi-choice format. A concrete example of the data creation process is shown in Figure 2. The benchmark comprises 6,027 aggregative questions derived from 182,330 real-world user-chatbot conversations, reflecting genuine user interests and societal trends, thus providing a resource for evaluating models’ ability to reason holistically at scale.
We evaluate current methods, including both non-reasoning and reasoning models, adapted to this task via fine-tuning, retrieval-augmented generation (RAG), and a customized retrieval approach developed specifically for aggregative reasoning: PROBE (Probing Retrieval Of Broad Evidence). Experimental results reveal substantial limitations in existing methods: current systems either struggle to reason effectively at scale or incur prohibitive computational costs. Even under idealized conditions—when the exact oracle contexts relevant to a query are provided—significant room for improvement remains; under realistic conditions, the performance further deteriorates.
Our findings emphasize the critical need for more scalable and effective methods capable of extracting collective insights from large-scale conversational datasets. While Aggregative Question Answering opens promising avenues for impactful real-world analytics, we acknowledge potential societal impacts, particularly when insights relate to sensitive topics like elections, public opinion, or public health—areas that could potentially be susceptible to manipulation. Nevertheless, we strongly believe that transparent, open academic research fosters responsible development and deployment of such powerful technologies. By introducing Aggregative Question Answering as a new task, we aim to spur future methods that fully harness the potential of large-scale conversational data, ultimately enabling deeper societal understanding and more responsive, informed applications of LLMs.
Our benchmark, code, and dataset are publicly available at https://github.com/yuntian-group/wildchat_aqa, and we also provide a user-friendly benchmark visualization tool at http://65.108.32.135:3000/dataview.
2 WildChat-AQA
To support research on Aggregative Question Answering, we constructed the WildChat-AQA benchmark based on the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat provides real-world conversations between users and chatbots, along with basic metadata such as timestamps and user locations. In this work, we extended these attributes by introducing additional attributes such as topics and keywords inferred from the conversation text using LLMs. These inferred attributes serve as the ground-truth annotations for constructing our benchmark. At evaluation time, models must infer them from conversations to answer aggregative questions. Table 1 summarizes the attributes and indicates which ones require inference and which ones are directly available.
| Name | Multi-Val | Inferred | Example |
| Location | No | No | United States, Canada |
| User Name | No | No | lostclasp37, toughcue8 |
| Time | No | No | 4/26/2023, 1:47:24 PM |
| Language | No | No | English, Russian |
| Topic | Yes | Yes | Software, Programming and Computer Science |
| Subtopic | Yes | Yes | Mobile Development, AI and ML |
| Keywords | Yes | Yes | C++, Pokémon |
2.1 Dataset Construction
The construction of WildChat-AQA involved four main steps, as illustrated in Figure 2:
Step 1: Preprocessing
We began by performing minHash-based deduplication (Hugging Face, 2023) to remove highly similar conversations, ensuring diversity. We also filtered conversations exceeding 4,096 tokens to maintain manageable context lengths. Additionally, we retained only active users (those with at least 10 interactions) to ensure sufficient user-specific data. We also generated user IDs from IP addresses and headers.
Step 2: Topic Discovery
To support meaningful aggregative queries, we prompted GPT-4o to summarize each conversation and extract relevant keywords. Using these summaries, we recursively applied TnT-LLM (Wan et al., 2024) to infer hierarchical topics at two levels: coarse-grained topics and fine-grained subtopics. Detailed prompts and examples can be found in Appendix E.
Step 3: Keywords Categorization
Certain subtopics, such as “Programming” and “Fan-fiction and Crossover,” contained many conversations. To support finer-grained aggregative queries, we further categorized keywords inferred from conversations into higher-level categories using LLMs so that we can derive aggregative information. For example, different Pokémon-related keywords (versions, characters, trademarks) were grouped into a single category “Pokémon”. Full details of this procedure are also available in Appendix E.
Step 4: Question Generation
Finally, we generated aggregative questions using combinations of attributes stored in our constructed database. We systematically sampled attribute combinations (from zero to three attributes as conditions) to query our database for a remaining target attribute. To ensure questions were both diverse and answerable, attribute combinations were carefully sampled. These structured queries were then translated into natural language using GPT-4.1.
2.2 Dataset Statistics
The resulting WildChat-AQA benchmark contains 182,330 user-chatbot conversations and 6,027 aggregative questions. These conversations cover 28 high-level topics, 455 fine-grained subtopics, and 14,482 keyword categories. Detailed statistics of questions organized by different attribute conditions and target attributes are shown in Table 8 in Appendix B. Unlike typical question-answering tasks, which derive answers from one or a few documents, WildChat-AQA requires models to reason over contexts whose total token counts range widely from to tokens. Figure 4 illustrates the distribution of context token counts. Full data statistics are provided in Appendix B.
2.3 Evaluation Protocol
We frame the evaluation of aggregative question answering as a ranking problem. During training, the model or system under evaluation is provided access to the entire WildChat-AQA dataset. At test time, the model is given an aggregative question along with 10 candidate answers. Its task is to rank these candidates according to their relevance to the question. We use standard ranking metrics NDCG@1, NDCG@3, NDCG@5, and NDCG@10 to measure performance.
2.4 Human Evaluation
To evaluate the quality of our inferred attributes, we conducted a human evaluation measuring both inter-annotator agreement (human-human) and human-model agreement using Cohen’s . Specifically, we randomly sampled 100 examples each for level-1 (topic) and level-2 (subtopic) taxonomy labeling. Due to the multi-label nature of these tasks, we computed per-label agreement by treating each possible category as an independent binary labeling task. For subtopic evaluation, we additionally reported macro-average agreement scores aggregated across all topics to provide a comprehensive view of annotation reliability.
| Name | Human–Human | Human-Model |
| Topic | 0.581 | 0.617 |
| Subtopic | 0.576 | 0.609 |
We found that Cohen’s for both topics and subtopics indicates moderate to substantial agreement (Cohen, 1960), demonstrating a high degree of reliability between human annotations and model predictions.
3 Probing Retrieval Of Broad Evidence
Traditional retrieval methods, including those used in retrieval-augmented generation (RAG), typically aim to identify a small set of highly specific, relevant documents. However, for Aggregative Question Answering, it is essential to identify a broader range of documents that collectively support reasoning about high-level aggregational insights. To address this unique requirement, we introduce a customized retrieval method, Probing Retrieval Of Broad Evidence (PROBE). PROBE operates in two main steps:
Broad Query Generation
Given a question , we first prompt a large language model to generate a comprehensive set of short, diverse queries that may help retrieve a broad range of relevant documents. Specifically, the LLM generates a set of queries related to the question. Additionally, the model generates strict filtering conditions to exclude documents clearly unrelated to the question. Formally, this process is defined as:
where represents the prompt.
Evidence Aggregation and Generation
Next, each generated query (alongside the filtering conditions ) is used individually to retrieve relevant documents. This results in separate retrieval runs. We then aggregate these results by merging the retrieved document lists according to their retrieval relevance scores. If a document appears multiple times across different queries, we use max pooling to assign it the highest relevance score it received from any query. Finally, we select the top documents from this aggregated list as evidence.
The resulting set of retrieved documents serves as supporting evidence for the model to perform aggregational reasoning and answer the question. An overview of the full PROBE retrieval pipeline is illustrated in Figure 3.
4 Experiments
We evaluate WildChat-AQA using various models, retrieval approaches, and conversation formats. We also explore model behaviors by performing ablation studies.
4.1 Models
4.2 Experimental Setups
We explore several experimental setups to investigate how effectively models leverage conversational data to answer aggregative questions:
No Context
The model directly answers questions without external inputs, relying solely on internal knowledge. This approach establishes baseline performance using only pre-existing knowledge. Due to resource constraints, we only evaluated this baseline using the strongest reasoning model (o4-mini).
Retrieval Augmented Generation (RAG)
We use standard retrieval-augmented generation using OpenAI’s text-embedding-3-large embeddings to retrieve relevant conversations as context.
Finetuning
We finetune pretrained models on the entire WildChat-AQA dataset.
PROBE
We use our proposed retrieval method, PROBE. Query generation uses GPT-4.1-mini, and retrieval relies on embeddings from OpenAI’s text-embedding-3-large model.
4.3 Raw vs. Summarized Document
Raw conversations are detailed but noisy (average 1,143.4 tokens each), whereas summarized conversations are more concise (average 21.5 tokens). Thus, we experimented with both raw and summarized conversation inputs to investigate their effectiveness for aggregative question answering. Implementation details for experiments are provided in Appendix D.
4.4 Main Results
| Model Name | Approach | Type | NDCG@1 | NDCG@3 | NDCG@5 | NDCG@10 | # Input Token (Million) |
| Random | None | / | 0.2501 | 0.3516 | 0.4368 | 0.6211 | - |
| o4-mini | None | / | 0.3063 | 0.4017 | 0.4805 | 0.6488 | 0.87 |
| Qwen3 8B | Finetune | Raw | 0.2694 | 0.3739 | 0.4589 | 0.6346 | 1.74 |
| Summary | 0.2984 | 0.3966 | 0.4807 | 0.6480 | 1.74 | ||
| Gemma3 4B | RAG | Raw | 0.3291 | 0.4356 | 0.5159 | 0.6688 | 73.48 |
| Summary | 0.3740 | 0.4895 | 0.5627 | 0.6991 | 174.62 | ||
| PROBE | Raw | 0.4766 | 0.5891 | 0.6478 | 0.7620 | 38.44 | |
| Summary | 0.5430 | 0.6513 | 0.6994 | 0.7969 | 17.35 | ||
| Qwen3 8B Think | RAG | Raw | 0.4168 | 0.5090 | 0.5779 | 0.7123 | 362.16 |
| Summary | 0.5273 | 0.6110 | 0.6646 | 0.7717 | 176.88 | ||
| PROBE | Raw | 0.6545 | 0.7305 | 0.7728 | 0.8483 | 315.52 | |
| Summary | 0.6944 | 0.7638 | 0.8005 | 0.8660 | 123.04 | ||
| Qwen3 32B Think | RAG | Raw | 0.4052 | 0.5020 | 0.5705 | 0.7081 | 182.90 |
| Summary | 0.5496 | 0.6321 | 0.6847 | 0.7850 | 176.88 | ||
| PROBE | Raw | 0.6525 | 0.7347 | 0.7759 | 0.8501 | 315.52 | |
| Summary | 0.7056 | 0.7753 | 0.8114 | 0.8725 | 123.04 | ||
| GPT-4.1 mini | RAG | Raw | 0.4494 | 0.5387 | 0.6035 | 0.7299 | 344.37 |
| Summary | 0.5782 | 0.6620 | 0.7104 | 0.8019 | 154.31 | ||
| PROBE | Raw | 0.6806 | 0.7536 | 0.7936 | 0.8628 | 298.69 | |
| Summary | 0.7308 | 0.7942 | 0.8282 | 0.8843 | 107.11 | ||
| o4-mini | RAG | Raw | 0.4730 | 0.5510 | 0.6116 | 0.7383 | 344.37 |
| Summary | 0.6122 | 0.6792 | 0.7242 | 0.8140 | 154.31 | ||
| PROBE | Raw | 0.7117 | 0.7747 | 0.8086 | 0.8745 | 298.69 | |
| Summary | 0.7571 | 0.8095 | 0.8386 | 0.8930 | 107.11 |
Table 3 presents performance results across different models, retrieval methods, and conversation formats.
Stronger models perform better.
Among tested models, o4-mini consistently achieved the highest performance, with a maximum NDCG@1 score of 0.7571. GPT-4.1-mini, while also strong, trailed slightly behind. Among open-source models, Qwen3-32B-think achieves the highest performance (0.7056 NDCG@1).
PROBE improves significantly over standard RAG.
Compared to standard RAG, PROBE consistently showed large performance improvements. On raw data, PROBE improved NDCG@1 scores by 14.8, 23.7, 24.7, 23.1, and 23.8 points for Gemma3-4B, Qwen3-8B-think, Qwen3-32B-think, GPT-4.1-mini, and o4-mini, respectively. A similar trend was observed using summarized conversations.
Summaries outperform raw conversations.
Models consistently performed better with summarized inputs, showing improved NDCG@1 scores of 4.5 to 14.4 points over raw conversations for standard RAG, and 4.0 to 6.6 points for PROBE. Summaries enable more efficient information retrieval and easier aggregation of insights.
Finetuning doesn’t help.
Finetuning Qwen3-8B did not significantly exceed random-chance performance. This result suggests that current finetuning methods fail to internalize aggregative information effectively from conversational datasets.
High token consumption.
Achieving good performance on this task required models to consume a very large number of input tokens as shown in Table 3. This highlights a significant computational challenge and motivates future research to improve efficiency.
| Method | R@5 | R@10 | R@20 | R@50 | R@100 | R@200 | R@500 |
| RAG-Dense | 0.01 | 0.02 | 0.04 | 0.07 | 0.10 | 0.14 | 0.21 |
| PROBE-Dense | 0.07 | 0.13 | 0.23 | 0.35 | 0.43 | 0.50 | 0.58 |
| - filter only | 0.05 (-0.02) | 0.09 (-0.04) | 0.16 (-0.07) | 0.24 (-0.11) | 0.29 (-0.14) | 0.33 (-0.17) | 0.40 (-0.18) |
| - question & filter | 0.06 (-0.01) | 0.12 (-0.01) | 0.21 (-0.02) | 0.32 (-0.03) | 0.40 (-0.03) | 0.46 (-0.04) | 0.53 (-0.05) |
| # Conversation | Approach | Recall | NDCG@5 |
| 5 | RAG | 0.01 | 0.5373 |
| PROBE | 0.07 | 0.6991 | |
[…]
| 20 | RAG | 0.04 | 0.5897 |
| PROBE | 0.23 | 0.7624 | |
[…]
| 50 | RAG | 0.07 | 0.6318 |
| PROBE | 0.35 | 0.7927 | |
[…]
| 200 | RAG | 0.14 | 0.6858 |
| PROBE | 0.50 | 0.8202 | |
[…]
| 500 | RAG | 0.20 | 0.7141 |
| PROBE | 0.58 | 0.8263 | |
[…]
4.5 Ablation Studies
We conducted ablation studies on a stratified 10% subset of the benchmark, selected based on condition and target types.
Retrieval Effectiveness is Critical
Retrieval performance has a significant impact on the final results. Table 5 presents the performance of o4-mini under varying recall rates achieved by different retrieval methods. We observe that higher recall rates consistently lead to improved NDCG scores.
Retrieval performance.
We compared various retrieval approaches, including vector-based embeddings, BM25, random, and ground-truth retrieval. Figure 5 shows recall rates for different retrieval strategies. PROBE consistently provided substantial improvements over standard RAG, with the highest recall from PROBE-Dense (summarized). Removing either the generated query or filtering steps notably degraded PROBE’s retrieval effectiveness (Table 4).
Current models lack effective aggregational reasoning
To explore model capabilities under ideal conditions, we performed experiments using oracle (ground-truth) documents as context (Table 6). All models performed better when given summarized contexts rather than raw conversations, indicating challenges in aggregating information from longer, noisier texts.
| Model Name | Ctx Type | NDCG@1 | NDCG@3 | NDCG@5 | NDCG@10 |
| Gemma3 4B | Raw | 0.4815 | 0.6057 | 0.6601 | 0.7703 |
| Summary | 0.5699 | 0.6787 | 0.7235 | 0.8102 | |
| Qwen3 8B Think | Raw | 0.7359 | 0.7991 | 0.8360 | 0.8894 |
| Summary | 0.7757 | 0.8268 | 0.8510 | 0.9003 | |
| Qwen3 32B Think | Raw | 0.7225 | 0.8044 | 0.8355 | 0.8897 |
| Summary | 0.8134 | 0.8605 | 0.8817 | 0.9199 | |
| GPT-4.1-mini | Raw | 0.7849 | 0.8388 | 0.8667 | 0.9121 |
| Summary | 0.8130 | 0.8602 | 0.8816 | 0.9216 | |
| o4-mini | Raw | 0.8003 | 0.8456 | 0.8719 | 0.9185 |
| Summary | 0.8478 | 0.8793 | 0.9005 | 0.9347 |
We further analyzed how performance varied with the number of provided conversations (Figure 6). Weaker models such as Gemma3 and Qwen3 showed a substantial performance gap between raw and summarized contexts, even when given the same number of conversations, highlighting their limited ability to implicitly extract relevant information. Stronger models like GPT-4.1-mini and o4-mini showed a smaller initial gap, but this gap widened notably when the context increased to 100 documents, demonstrating that even advanced models struggle with aggregating and reasoning effectively over extensive raw contexts.
Performance improves with more context.
Unlike standard RAG tasks, Aggregative Question Answering fundamentally relies on a broader set of documents. As more documents were provided, models improved significantly in answering aggregative questions (Figure 7). This finding validates that aggregative question answering requires extensive context and global dataset knowledge.
Aggregative question answering is reasoning-intensive.
We evaluated Qwen3-32B (with optional reasoning via “think” mode) to measure the effect of explicit reasoning. Results (Table 7) consistently showed reasoning led to significant performance improvements across all experimental setups, indicating aggregative question answering demands substantial reasoning abilities.
| Method | NDCG@1 | NDCG@3 | NDCG@5 | NDCG@10 |
| Oracle | 0.72 | 0.79 | 0.82 | 0.88 |
| + thinking | 0.81 (+0.09) | 0.86 (+0.07) | 0.88 (+0.06) | 0.92 (+0.04) |
| RAG (Summary) | 0.48 | 0.56 | 0.62 | 0.74 |
| + thinking | 0.54 (+0.07) | 0.62 (+0.07) | 0.68 (+0.06) | 0.78 (+0.04) |
| PROBE (Summary) | 0.64 | 0.71 | 0.75 | 0.84 |
| + thinking | 0.68 (+0.04) | 0.76 (+0.05) | 0.80 (+0.05) | 0.86 (+0.02) |
5 Future Research Directions
Reasoning Over Very Long Context
In this work, we experimented with several reasoning-capable models and observed that current models typically have limited context windows, and performance degrades sharply as the length of the input context increases. Developing efficient and accurate methods for reasoning over very long textual contexts remains an important open problem.
Cost-Efficient Aggregative Question Answering
Current effective solutions for Aggregative Question Answering require processing extremely large amounts of text, resulting in substantial computational costs. Future research could explore hierarchical indexing, retrieval strategies, and long-term memory mechanisms to reduce token consumption and improve computational efficiency.
Streaming Aggregative Question Answering
In real-world scenarios, chatbot conversations often arrive in continuous streams rather than static collections. Future research could explore methods to dynamically update aggregational insights as new interactions occur in real time. Ideally, conversational agents would continuously integrate information from ongoing interactions—similar to how humans update their understanding based on new experiences—to maintain up-to-date and adaptive aggregational knowledge.
6 Conclusion
In this paper, we introduced Aggregative Question Answering, a new task aimed at extracting collective insights from large-scale conversational data generated by interactions between users and LLM-powered chatbots. To facilitate research in this area, we constructed the WildChat-AQA benchmark, comprising 6,027 aggregational questions derived from 182,330 real-world chatbot conversations. Our experiments demonstrate that existing state-of-the-art methods, including fine-tuning, retrieval-augmented generation (RAG), and even an improved RAG approach specifically adapted for this task—PROBE (Probing Retrieval Of Broad Evidence)—struggle significantly, either failing to reason effectively at the necessary global scale or incurring prohibitively high computational costs. Looking ahead, we believe addressing these challenges would enable future models to better derive meaningful user and societal insights from large-scale conversational data.
Limitations
Potential Errors in Model-derived Annotations
Although we employed powerful large language models and pipelines (such as GPT-4o and TnT-LLM) to infer attributes and assign taxonomy labels, errors and inconsistencies may still occur due to model hallucinations or instruction misalignment. Additionally, real-world conversational data is inherently noisy, ambiguous, and challenging to categorize neatly, making it impossible to ensure completely error-free annotations.
Artificiality of Generated Questions
Aggregative questions in WildChat-AQA were generated by prompting GPT-4.1 to translate structured database queries into natural-language questions. While effective, this method may result in questions that still feel somewhat templated or artificial, potentially limiting their naturalness and realism compared to genuinely human-authored queries.
Ethical Considerations
Aggregative Question Answering opens promising avenues for impactful real-world analytics but also raises potential ethical and societal concerns, particularly when insights relate to sensitive topics such as elections, public opinion, or public health—areas that could potentially be susceptible to manipulation. To mitigate such risks, all experiments conducted in this work rely exclusively on the publicly available and anonymized WildChat dataset, which is explicitly intended for open research purposes (licensed under ODC-BY). By introducing WildChat-AQA as an open benchmark, we aim to empower transparent academic research that responsibly explores both the capabilities and risks associated with aggregational analytics. Our goal is to encourage the open research community to better understand and evaluate these powerful systems, rather than relying solely on proprietary analyses conducted behind closed doors.
Appendix A Related Works
Question Answering
Question answering typically involves a diverse range of perspectives. Datasets such as TriviaQA Joshi et al. (2017), RACE Lai et al. (2017), HotPotQA Yang et al. (2018), Natural Questions Kwiatkowski et al. (2019), MuSiQue Trivedi et al. (2022), 2Wiki Ho et al. (2020), PopQA Mallen et al. (2023), and MultiHop-RAG Tang and Yang (2024) focus on local information, where answers can be derived from one or several documents. In contrast, other benchmarks such as MMLU Hendrycks et al. (2021a), MATH Hendrycks et al. (2021b), GSM8K Cobbe et al. (2021), and Big-Bench bench authors (2023) emphasize science, technology, engineering, mathematics, and logical reasoning. These primarily evaluate models’ world knowledge and reasoning capabilities but lack a benchmark for understanding large-scale datasets and deriving high-level insights. Recent works such as GraphRAG Edge et al. (2025) address the long-context challenge by extracting entities and relationships from extended text data and constructing graph structures to answer questions.
Long Context Retrieval Augmented Generation
Lewis et al. (2020) has emerged as a prominent approach for enhancing the performance of large language models (LLMs) on knowledge-intensive tasks while also mitigating hallucinations. Recently, advances in computational capabilities have spurred interest in extending RAG to support very long contexts. Several studies—such as those by Jiang et al. (2024), Zhao et al. (2024a), and Jin et al. (2025)—have proposed methods to improve the effectiveness of LLMs in long-context settings. In parallel, Lee et al. (2024) introduced LOFT, a new benchmark designed to evaluate LLMs on a broad range of tasks addressable by either RAG or long-context modeling.
Summarization
Summarization has been a long-standing challenge in natural language processing. Early benchmark datasets, such as CNN/Daily Mail See et al. (2017) and XSum Narayan et al. (2018), primarily targeted single-document summarization. Subsequent efforts, including MultiNews Fabbri et al. (2019) and MS2 DeYoung et al. (2021), extended this task to the multi-document setting. Another line of related work focuses on query-based summarization, for which QMSum Zhong et al. (2021) and DUC 2005 Dang (2006) are two widely used datasets.
Text to SQL
Text-to-SQL is a widely studied approach for tackling aggregative question answering. In this paradigm, the model is required to generate a structured database query based on a natural language question. Several established benchmarks have been proposed to evaluate this task, including WikiSQL Zhong et al. (2017), Spider Lei et al. (2024), BIRD Li et al. (2024b), and WikiTableQA Pasupat and Liang (2015). Additionally, LOFT Lee et al. (2024) includes a sub-task specifically designed to assess how effectively large language models can emulate database-style querying.
Appendix B Data Statistics
B.1 Statistics of Generated Question by Condition and Targets
| Condition | Target | Count |
| 0 Condition | ||
| none | topic | 1 |
| none | loc | 1 |
| none | lang | 1 |
| 1 Condition | ||
| user | keywords | 370 |
| user | time | 100 |
| keywords | user | 96 |
| user | lang | 60 |
| user | topic | 54 |
| time | user | 39 |
| topic | subtopic | 26 |
| loc | topic | 20 |
| loc | keywords | 17 |
| lang | topic | 9 |
| time | topic | 6 |
| time | keywords | 6 |
| topic | loc | 6 |
| topic | user | 6 |
| topic | lang | 4 |
| topic | keywords | 4 |
| time | lang | 4 |
| lang | keywords | 1 |
| 2 Conditions | ||
| user, topic | subtopic | 199 |
| user, topic | keywords | 185 |
| user, user | subtopic | 141 |
| user, topic | time | 114 |
| topic, lang | subtopic | 100 |
| time, topic | user | 98 |
| time, topic | subtopic | 98 |
| topic, lang | user | 98 |
| topic, loc | time | 97 |
| topic, keywords | user | 97 |
| Condition | Target | Count |
| topic, loc | subtopic | 96 |
| topic, keywords | time | 96 |
| time, user | keywords | 94 |
| topic, subtopic | user | 93 |
| subtopic, subtopic | user | 93 |
| topic, loc | keywords | 82 |
| topic, lang | time | 74 |
| time, topic | loc | 60 |
| topic, subtopic | keywords | 55 |
| topic, topic | user | 55 |
| time, user | topic | 53 |
| user, user | topic | 53 |
| time, topic | keywords | 49 |
| topic, subtopic | loc | 39 |
| time, loc | topic | 34 |
| time, lang | topic | 31 |
| topic, lang | keywords | 27 |
| time, topic | lang | 15 |
| topic, subtopic | lang | 13 |
| topic, loc | user | 10 |
| 3 Conditions | ||
| loc, topic, subtopic | user | 287 |
| lang, topic, subtopic | user | 284 |
| user, topic, subtopic | keywords | 276 |
| time, loc, topic | user | 199 |
| time, topic, subtopic | keywords | 175 |
| user, user, user | subtopic | 132 |
| user, topic, keywords | time | 114 |
| time, topic, keywords | user | 100 |
| time, loc, topic | subtopic | 100 |
| time, user, topic | subtopic | 100 |
| loc, topic, keywords | user | 99 |
| user, topic, subtopic | time | 98 |
| user, topic, keywords | subtopic | 98 |
| loc, topic, keywords | time | 98 |
| lang, topic, keywords | time | 98 |
| time, topic, subtopic | user | 97 |
| lang, topic, keywords | user | 96 |
| topic, subtopic, keywords | user | 94 |
| loc, topic, subtopic | keywords | 93 |
| lang, topic, subtopic | keywords | 82 |
| time, topic, subtopic | loc | 76 |
| user, user, user | topic | 51 |
B.2 Language Distribution
We provide a statistics of all language involved in the conversations in Table 9.
| Language | Count | Language | Count | Language | Count | Language | Count |
| English | 124,646 | Spanish | 4,193 | Italian | 744 | Polish | 527 |
| Russian | 22,877 | Portuguese | 3,532 | Korean | 605 | Vietnamese | 463 |
| Chinese | 6,434 | Turkish | 1,408 | Indonesian | 566 | Ukrainian | 406 |
| French | 4,782 | Latin | 1,239 | Dutch | 549 | Other | 1,824 |
| German | 4,487 | Arabic | 863 | Tagalog | 537 |
B.3 Keywords Cloud
To illustrate the result of keywords categorization, we build a keywords cloud in Figure 8
B.4 Topic and Subtopic overview
| Parent Topic | Sub-topic | Count |
| Creative Writing and Fiction | Dialogue & Scripted Scenes | 25421 |
| Fanfiction & Universe Crossovers | 20323 | |
| Extended Narrative Prose | 19771 | |
| Humorous & Satirical Narratives | 11901 | |
| Erotic & Sensual Narratives | 8304 | |
| World-Building & Adventure Narratives | 6470 | |
| Creative Naming & Prompt Generation | 4388 | |
| Sports & Competition Narratives | 3370 | |
| Transformation & Identity Narratives | 3283 | |
| Character Profiles & Descriptions | 2025 | |
| Fictional News & Media Formats | 1912 | |
| Poetic & Lyric Composition | 1608 | |
| Interactive & Roleplaying Narratives | 827 | |
| Law, Regulation and Criminal Justice | Violent Crimes | 630 |
| Regulatory Compliance and Licensing | 454 | |
| Civil Litigation and Consumer Protection | 284 | |
| Employment and Labor Law | 198 | |
| Sexual Crimes | 183 | |
| Intellectual Property and Copyright | 163 | |
| Financial, Fraud, and Cyber Offenses | 142 | |
| Robbery, Theft, and Property Offenses | 130 | |
| Judicial Process and Court Administration | 117 | |
| Constitutional Rights and Civil Liberties | 81 | |
| Terrorism, War Crimes, Treason, and Political Violence | 68 | |
| Corruption and Abuse of Power | 64 | |
| Public Order Offenses | 54 | |
| Immigration and Border Control | 51 | |
| Drug-Related Offenses | 50 | |
| Family and Marital Law | 48 | |
| Entertainment, Games, and Media | Fanfiction & Crossovers | 25629 |
| Original Fiction & Scripts | 4834 | |
| NSFW & Explicit Scenes | 3717 | |
| Live-Action Film & TV | 2963 | |
| Western Animation & Comics | 2048 | |
| Gaming Story & Lore | 1895 | |
| Celebrity & Pop Culture | 1882 | |
| Gaming Mechanics & Tech | 1660 | |
| Music & Stage | 1651 | |
| Sports, eSports, & Pro Wrestling | 1557 | |
| Anime & Manga | 1552 | |
| Production & Broadcasting | 1044 | |
| Tabletop & TTRPG | 804 | |
| Software, Programming and Computer Science | Programming | 17413 |
| Web Development | 3603 | |
| AI and Machine Learning | 2787 | |
| Cybersecurity | 1930 | |
| Game Development, Design, and Modding | 1737 | |
| Databases and Queries | 1724 | |
| Operating Systems and Administration | 1414 | |
| Productivity and Desktop Software | 1215 | |
| Computer Networking | 1176 | |
| DevOps and Cloud | 1083 | |
| Data Analysis, Visualization and Business Intelligence | 1031 | |
| Mobile Development and Mobile Apps | 972 | |
| Computer Graphics | 740 | |
| Computer Science Theory | 612 | |
| Computer Hardware, Architecture, and Peripherals | 576 | |
| Software Architecture and Software System Design | 438 | |
| Testing and Quality Assurance | 350 | |
| Blockchain and Cryptocurrency | 336 | |
| Embedding Systems and IoT | 286 | |
| Human Computer Interaction | 184 | |
| Software Development Methodology and Project Management | 165 | |
| Science, Mathematics and Logical Reasoning | Physics: Mechanics, Thermodynamics, and Fields | 1877 |
| Basic Arithmetic and Numbers | 1376 | |
| Organismal Biology and Evolution | 1360 | |
| General Chemistry and Reactions | 1339 | |
| Cellular and Medical Sciences | 1239 | |
| Astronomy and Astrophysics | 1130 | |
| Earth Science and Environment | 1031 | |
| Statistics and Probability | 912 | |
| Algebra and Vectors | 833 | |
| Logic and Puzzles | 795 | |
| Geometry and Trigonometry | 724 | |
| Computational Science and Modeling | 610 | |
| Calculus and Higher Mathematics | 505 | |
| Materials, Engineering, and Technology | 363 | |
| Personal Advice and Support | Navigating Romance and Dating | 464 |
| Enhancing Personal Growth and Discipline | 286 | |
| Building Communication and Social Skills | 164 | |
| Offering Emotional Support and Love | 137 | |
| Navigating Sexual Intimacy, Consent, and Well-Being | 128 | |
| Supporting Mental Health and Well-Being | 111 | |
| Guiding Family, Parenting, and Caregiving | 99 | |
| Boosting Self-Confidence and Esteem | 81 | |
| Handling Career and Workplace Challenges | 73 | |
| Exploring Personal Values and Choices | 70 | |
| Seeking Apologies, Forgiveness, and Trust | 65 | |
| Addressing Financial Management and Housing | 47 | |
| Improving Physical Health and Body Image | 47 | |
| Managing Unwanted Contact and Boundaries | 38 | |
| Seeking Legal Guidance and Protective Measures | 34 | |
| Embracing Identity and Lifestyle Transitions | 32 | |
| Recovering from Breakups and Heartache | 32 | |
| Handling Emergencies, Threats, or Crises | 30 | |
| Overcoming Addictions and Harmful Habits | 19 | |
| Coping with Grief and Loss | 15 | |
| Business, Commerce and Finance | Digital Marketing & Social Media | 4010 |
| Investments & Financial Markets | 934 | |
| Business Operations & Quality Management | 914 | |
| Accounting & Financial Reporting | 891 | |
| Economic Trends & Macro Outlook | 739 | |
| Corporate Governance & Leadership | 492 | |
| Customer Service & Complaints | 460 | |
| Legal & Regulatory Compliance | 435 | |
| Supply Chain & Logistics | 426 | |
| Wholesale & B2B Distribution | 404 | |
| Banking & Monetary Policies | 402 | |
| Careers & Professional Development | 373 | |
| Entrepreneurship & Startups | 356 | |
| History and Culture | Modern and Contemporary History (19th Century–Present) | 1407 |
| Conflicts and Wars | 1088 | |
| Medieval Europe | 716 | |
| Philosophy and Political Ideologies | 624 | |
| Art, Architecture, and Heritage | 616 | |
| Religion and Theology | 513 | |
| Traditions, Customs, and Rituals | 395 | |
| Popular Culture and Mass Media | 388 | |
| Pre-Modern East Asia | 386 | |
| Colonialism, Imperialism, and Independence | 343 | |
| Ancient Non-Classical Civilizations | 322 | |
| Classical Rome | 269 | |
| Diplomacy and Treaties | 251 | |
| Language and Literature | 240 | |
| Archaeology and Ancient Technologies | 217 | |
| Sports and Leisure | 197 | |
| Civil Rights and Social Justice | 192 | |
| Ancient Greece and Hellenic Culture | 174 | |
| Legal Systems and Codes | 172 | |
| Social Hierarchies and Slavery | 170 | |
| Myths and Folklore | 166 | |
| Gender and Women’s History | 166 | |
| Indigenous Peoples | 157 | |
| Science and Medicine | 154 | |
| Islamic and Middle Eastern Empires | 119 | |
| Exploration and Discoveries | 100 | |
| Lifestyle and Hobbies | Exploring fashion and accessories | 204 |
| Hair and Personal Grooming | 189 | |
| Beauty, makeup, and self-care | 110 | |
| Health, sports, and active living | 107 | |
| Minimalist living and conscious habits | 95 | |
| Personal expression, identity, and body positivity | 81 | |
| Creative crafts and DIY projects | 67 | |
| Outdoor Recreation and Camping | 61 | |
| Relationships, family, and social bonding | 59 | |
| Pets, animals, and responsible care | 46 | |
| Spirituality, meditation, and mindfulness | 45 | |
| Music, dance, and performing arts | 43 | |
| Games, collecting, and playful hobbies | 42 | |
| Social events, parties, and gatherings | 40 | |
| Costumes and cosplay | 37 | |
| Cooking, baking, and culinary hobbies | 31 | |
| Productivity and time management | 30 | |
| Travel, tourism, and new adventures | 24 | |
| Digital lifestyle and social media presence | 24 | |
| Seasonal festivities and holiday decorating | 12 | |
| Gardening and horticulture | 7 | |
| Home organization and interior comfort | 6 | |
| Academic Resource, Education and Learning | Academic Research, Methods, and Presentation | 801 |
| Curriculum and Course Development | 697 | |
| STEM and Technical Education | 428 | |
| Teaching Strategies and Pedagogical Tools | 423 | |
| Health and Medical Education | 326 | |
| Technology and AI Integration in Education | 296 | |
| Professional and Vocational Training | 248 | |
| Educational Policy and Leadership | 195 | |
| University Admissions and Scholarship Guidance | 157 | |
| Language Learning and Translation | 135 | |
| Memory, Study, and Exam Strategies | 118 | |
| Creative Arts and Literature in Education | 110 | |
| Early Childhood Education and Development | 104 | |
| Special Education and Inclusive Learning | 66 | |
| Socio-Emotional Learning and Wellbeing | 60 | |
| Environmental and Social Education | 43 | |
| Academic Ethics and Publication Guidelines | 34 | |
| Parental Engagement and Child Education | 34 | |
| Classroom Management and Student Engagement | 25 | |
| Undefined | 2 | |
| Psychology, Mental Health and Emotional Support | Communication Skills & Empathy | 211 |
| Child & Adolescent Mental Health | 199 | |
| Relationship & Interpersonal Challenges | 181 | |
| Stress, Coping Strategies & Resilience | 158 | |
| Mood Disorders (Depression & Bipolar) | 155 | |
| Anxiety, Panic & Phobias | 112 | |
| Psychological Theories & Historical Perspectives | 109 | |
| Therapy & Counseling Methods | 103 | |
| Sexual Orientation, Gender & Sexual Behaviors | 102 | |
| Trauma & PTSD | 99 | |
| Emotional Support for Crises & Suicidal Ideation | 97 | |
| Self-esteem & Self-sabotage | 95 | |
| Neurodevelopmental Disorders (ADHD, Autism, etc.) | 90 | |
| Addiction & Substance Use | 69 | |
| Abuse, Violence & Bullying | 67 | |
| Grief & Loss | 54 | |
| Personality Disorders | 42 | |
| Schizophrenia & Psychotic Symptoms | 38 | |
| Social & Cultural Factors in Mental Health | 37 | |
| Sleep & Dream Analysis | 36 | |
| Dissociative Disorders & Maladaptive Daydreaming | 33 | |
| Medication & Pharmacological Discussions | 28 | |
| Eating & Body Image Disorders | 25 | |
| Obsessive & Compulsive Disorders | 16 | |
| Interactive Activities with AI Chatbots | Explicit or Sexual Roleplay | 1023 |
| Developer Mode or Policy-Breaking Requests | 456 | |
| Interactive Storytelling with User Control | 380 | |
| Comedic or Vulgar Roleplay | 256 | |
| Flirty or Romantic Scenarios | 217 | |
| Childlike or Energetic Roleplay | 188 | |
| Game or Puzzle Interactions | 162 | |
| Roleplay with Personal or Close Relationships | 112 | |
| Fantasy or Mythical Adventures | 101 | |
| Roleplay with Non-Human Traits | 78 | |
| Action or Combat-Based Roleplay | 77 | |
| Testing Chatbot’s Memory or Logic | 68 | |
| Roleplay with Theatrical or Literary Flair | 60 | |
| Roleplay with Real-World Professions | 49 | |
| Minimalistic or Symbolic Responses Only | 44 | |
| Roleplay with Custom Machinery or System Simulation | 43 | |
| Roleplay with Worship or Devotion | 37 | |
| Roleplay with Social or Political Themes | 29 | |
| Roleplay as Rebels or Criminals | 27 | |
| Hypnosis or Therapeutic Roleplay | 7 | |
| Linguistics, Language and Translation | Rewriting and Paraphrasing | 8331 |
| Translation | 7997 | |
| Vocabulary and Terminology | 2586 | |
| Proof Reading and Grammar Correction | 2102 | |
| Linguistic Analysis | 1099 | |
| Summarization | 779 | |
| Language Learning Assistance | 503 | |
| Phonetics and Pronunciation | 464 | |
| Information Extraction | 391 | |
| Social Issues, Politics and Governance | Domestic Governance & Public Policy | 1334 |
| Political Theories & Ideological Debates | 1231 | |
| International Relations & Geopolitics | 1190 | |
| Social Justice, Identity & Cultural Norms | 1009 | |
| Political Leadership & Electoral Dynamics | 742 | |
| National Security & Crisis Management | 543 | |
| Economic Policy & Regulation | 366 | |
| Medicine and Health | Orthopedics and Musculoskeletal Health | 467 |
| Nutrition and Dietary Supplements | 466 | |
| Infectious Diseases and Vaccines | 385 | |
| Rehabilitation and Recovery | 384 | |
| Pharmacology and Medication Safety | 378 | |
| Eye, ENT, and Respiratory Conditions | 376 | |
| Surgery and Emergency Care | 341 | |
| Mental Health and Wellbeing | 328 | |
| Reproductive Health and Childbirth | 313 | |
| Digestive, Metabolic, and Endocrine Disorders | 304 | |
| Sexual Health and Function | 243 | |
| Healthcare Systems and Public Health | 238 | |
| Neurology and Nervous System Disorders | 212 | |
| Dermatology and Skin Care | 201 | |
| Diagnostic Tests and Imaging | 190 | |
| Cardiovascular Diseases and Hypertension | 181 | |
| Exercise, Fasting, and Weight Control | 177 | |
| Pediatrics and Child Health | 169 | |
| Preventive Medicine and Wellness | 152 | |
| Cancer and Oncological Care | 141 | |
| Medical Technology and Telemedicine | 109 | |
| Oral Health and Dentistry | 103 | |
| Substance Use and Addiction | 96 | |
| Allergies and Immune Conditions | 88 | |
| Occupational and Environmental Health | 80 | |
| Genetics and Rare Conditions | 76 | |
| Veterinary Medicine and Animal Health | 42 | |
| Technology, Engineering and Industry | Mechanical Engineering and Manufacturing | 678 |
| Electrical and Electronics Design | 418 | |
| Materials Science and Engineering | 405 | |
| Aerospace and Space Exploration | 381 | |
| Consumer Electronics and Gadgets | 364 | |
| Big Data, IoT, and Smart Systems | 310 | |
| Blockchain and Decentralized Tech | 305 | |
| Networking, Telecommunications, and Cybersecurity | 287 | |
| Civil Engineering and Infrastructure | 278 | |
| Automotive Engineering and Vehicle Technology | 257 | |
| AI and Machine Learning | 251 | |
| VR, AR, and XR Solutions | 245 | |
| Industrial Safety and Compliance | 220 | |
| Robotics, Drones, and Mechatronics | 203 | |
| Military and Defense Technology | 185 | |
| Energy and Sustainable Manufacturing | 156 | |
| Cloud, Virtualization, and Enterprise Platforms | 131 | |
| Supply Chain and Logistics Management | 115 | |
| Software Development and Web Frameworks | 108 | |
| Quantum and High-Performance Computing | 101 | |
| Agricultural Engineering and Food Industry | 84 | |
| Digital Media, Broadcasting, and Streaming | 75 | |
| Hardware Innovation and CPU/GPU Development | 68 | |
| HCI, UI/UX, and Interactive Tech | 67 | |
| Marine and Offshore Engineering | 62 | |
| Data Storage and Retention | 61 | |
| Engineering Education and STEM Training | 55 | |
| Biomedical, Biotech, and Wearables | 55 | |
| Gaming Technology and eSports | 46 | |
| Industrial Digitalization and Change Management | 37 | |
| Product Design and Industrial Innovation | 29 | |
| 3D Printing and Additive Manufacturing | 16 | |
| General Digital Support | AI Capabilities | 472 |
| AI Limitations | 397 | |
| AI Identity, Version, and Origins | 161 | |
| Correcting or Revising AI Responses | 61 | |
| Technical Guidance: External Apps and Websites | 57 | |
| AI Emotions or Opinions | 48 | |
| Creative Writing | 38 | |
| Official Links or Verification | 33 | |
| Coding Tasks | 29 | |
| Technical Guidance: Phones and Software | 24 | |
| Email and Account Management | 19 | |
| Comparison with Other AI Systems | 18 | |
| Education or Research Use | 17 | |
| Search and Browsing Advice | 10 | |
| Payment or Subscription | 5 | |
| Food, Cooking and Nutrition | Nutritional Guidance & Diet Planning | 569 |
| Recipes & Cooking Techniques | 518 | |
| Ingredient Selection & Quality | 218 | |
| Culinary Culture & Dining Experience | 166 | |
| Food Safety & Storage | 76 | |
| Art and Design | Product & Merchandise Design | 1086 |
| AI-Generated Art & Prompt Engineering | 585 | |
| Digital Media & Advertising Design | 492 | |
| Color Theory & Visual Composition | 407 | |
| Character & Animation Design | 290 | |
| Art History & Critique | 270 | |
| Editorial & Commercial Illustration | 262 | |
| Fashion & Costume Design | 252 | |
| Logo & Branding Design | 213 | |
| Educational & Children’s Art | 204 | |
| Architectural & Environmental Design | 192 | |
| Digital Art & Software Techniques | 132 | |
| Traditional & Manual Art Techniques | 116 | |
| Religion, Mythology and Spirituality | Biblical and Scriptural Narratives | 981 |
| Islamic Sacred Narratives | 363 | |
| Classical Mythology Narratives | 356 | |
| Eastern Sacred Narratives | 243 | |
| Modern Esoteric and Occult Spirituality | 188 | |
| Religion, Society, and Cultural Critique | 178 | |
| Astrological and Divinatory Traditions | 169 | |
| Folk and Indigenous Myth Narratives | 164 | |
| Norse and Germanic Mythological Narratives | 44 | |
| Ancient Near Eastern and Persian Narratives | 31 | |
| Literature and Book Analysis | Narrative and Prose Analysis | 1482 |
| Poetry and Versified Analysis | 427 | |
| Literary Guidance and Recommendations | 355 | |
| Advanced Literary Criticism | 43 | |
| Philosophy and Ethics | Epistemology, Logic, and Fallacies | 349 |
| Law, Governance, and Political Philosophy | 341 | |
| Mind, Consciousness, and Reality | 303 | |
| Religion, Theology, and Faith Traditions | 299 | |
| Existentialism, Death, and Meaning | 176 | |
| Moral Theories, Virtue, and Character Development | 171 | |
| Moral Speech and Expression | 146 | |
| Critical Theory and Postmodernism | 133 | |
| Consent, Power, and Manipulation | 104 | |
| Cultural Norms and Social Ethics | 100 | |
| Aesthetics and Artistic Philosophy | 91 | |
| Ethics in AI and Future Technologies | 90 | |
| Professional Ethics and Duty | 81 | |
| Markets, Capitalism, and Economic Fairness | 43 | |
| Bioethics, Medicine, and Life Origins | 42 | |
| Morality Toward Animals | 40 | |
| Love, Relationships, and Emotional Ethics | 28 | |
| Environmental Ethics and Sustainability | 19 | |
| Sports and Athletics | NCAA College Football | 1012 |
| Motorsport | 607 | |
| NBA Basketball | 604 | |
| NCAA College Basketball | 549 | |
| Global Soccer | 538 | |
| Fictional or Hypothetical Scenarios | 451 | |
| Professional American Football | 313 | |
| General or Cross-Sport Training & Fitness | 218 | |
| Professional Wrestling | 146 | |
| Baseball | 68 | |
| Combat Sports | 64 | |
| Cricket | 60 | |
| Cycling (Races & Gear) | 59 | |
| Ice Hockey | 25 | |
| Tennis and Other Racket Sports | 18 | |
| Rugby | 14 | |
| Gymnastics & Swimming | 7 | |
| Volleyball | 3 | |
| Golf | 2 | |
| Environment, Ecology and Sustainability | Climate Change Causes, Impacts, and Adaptation | 140 |
| Biodiversity Conservation and Wildlife Protection | 119 | |
| Greenhouse Gas Emissions and Carbon Management | 117 | |
| Pollution (Air, Water, Soil) and Remediation | 102 | |
| Waste Management and Circular Economy | 101 | |
| Environmental Policies, Laws, and Regulations | 82 | |
| Sustainable Energy and Energy Transition | 74 | |
| Green Industry, Corporate Sustainability, and Innovation | 72 | |
| Water Resource Management and Conservation | 67 | |
| Ecological Economics and Sustainable Development | 66 | |
| Environmental Education and Public Awareness | 45 | |
| Deforestation, Reforestation, and Sustainable Forestry | 43 | |
| Environmental Monitoring, Data Analysis, and Reporting | 40 | |
| Sustainable Lifestyles and Consumer Choices | 39 | |
| Sustainable Packaging, Recycling, and Plastics Reduction | 37 | |
| Sustainable Agriculture and Food Systems | 35 | |
| Marine and Coastal Conservation | 33 | |
| Sustainable Cities and Urban Development | 33 | |
| Ecological Restoration and Ecosystem Management | 33 | |
| Digital Technologies and Sustainability | 32 | |
| Sustainable Architecture and Construction | 26 | |
| Sustainable Transportation and Mobility | 23 | |
| Soil Health and Land Use Management | 22 | |
| Environmental Disaster Preparedness and Risk Reduction | 20 | |
| Carbon Markets and Climate Finance | 19 | |
| Eco-friendly Materials and Green Design | 17 | |
| Community-based Conservation and Participation | 15 | |
| Climate Negotiations and International Agreements | 12 | |
| Protected Areas and Natural Heritage Sites | 12 | |
| Environmental and Climate Justice | 11 | |
| Conservation Technology and Innovation | 6 | |
| Environmental Impact Assessment and Life Cycle Analysis | 5 | |
| Sustainable Tourism and Ecotourism | 3 | |
| Travel and Tourism | Cultural, Heritage & City Experiences | 126 |
| Transport & Logistics | 87 | |
| Travel Itineraries & Trip Planning | 65 | |
| Accommodation & Lodging | 54 | |
| Tourism Industry, Policy & Market | 49 | |
| Culinary & Dining | 40 | |
| Visa & Travel Documentation | 40 | |
| Beach, Coastal & Cruise Tourism | 37 | |
| Entertainment & Nightlife | 28 | |
| Adventure & Outdoor Activities | 25 | |
| Professional Development and Career Advice | Cover Letters & SOPs | 270 |
| Resume & CV Enhancement | 233 | |
| Workplace Culture & Dynamics | 132 | |
| Skill Development & Advanced Education | 128 | |
| Leadership & Team Management | 106 | |
| Salary & Compensation Guidance | 96 | |
| Recruitment & Talent Acquisition | 96 | |
| Industry-Specific Career Advice | 75 | |
| LinkedIn & Personal Branding | 69 | |
| Job Search & Networking Strategies | 60 | |
| Career Transitions & Upskilling | 60 | |
| Negotiation & Employment Contracts | 42 | |
| Interview Preparation & Techniques | 31 | |
| Employment Documentation & Verification | 31 | |
| Freelancing & Entrepreneurship | 19 | |
| Home and Household | Gardening: Planting & General Care | 140 |
| Gardening: Soil & Fertilization | 128 | |
| Fruit & Berry Cultivation | 107 | |
| Home Fixtures & Materials | 83 | |
| Gardening: Pest & Disease Management | 75 | |
| Interior Design & Decoration | 60 | |
| Home Maintenance & Appliance Repair | 54 | |
| Laundry & Fabric Care | 36 | |
| DIY Tools & Household Projects | 31 | |
| Household Cleaning & Stain Removal | 27 | |
| Outdoor Landscaping & Mulching | 24 | |
| Eco-Friendly & Sustainable Practices | 15 | |
| Household Safety & Security | 14 | |
| Real Estate & Tenancy | 13 | |
| Household Management & Lifestyle | 13 | |
| Home Organization & Storage Solutions | 8 | |
| Household Pets & Animal Care | 5 |
Appendix C Data Visualization Demonstration
We developed an interactive data visualization interface using React.js and Next.js for the frontend, and FastAPI for the backend implementation. MongoDB serves as the database system. An overview of the interface is shown in Figure 9. Users can filter generated questions using a configurable question filter, as illustrated in Figure 10.
The filtering mechanism allows users to select one or more attributes for both the condition and target fields to retrieve relevant questions. For instance, the filters “user_pair” and “user_triplet” refer to questions based on common interests between two or three users, respectively. Similarly, “joint_topic” and “joint_subtopic” denote filters that select conversations involving shared topics or subtopics.
For each question, the interface displays the number of supporting dialogues and their associated token counts. Additional distributions—such as raw keywords, aggregated keywords, language, topic, location, and user identity—are visualized to facilitate deeper insights.
Users can also explore the “DIALOGUES” panel to view all conversation excerpts that support a particular question. Each dialogue entry includes detailed metadata: username, timestamp, topic, subtopic, generated summary, raw extracted keywords, and aggregated keywords. This comprehensive display allows users to audit or explore the basis of each proposed question in context.
Appendix D Experiment Implementation Details
We employed MongoDB v8.0.4 for question proposal generation and ground-truth-based retrieval. All retrieval experiments utilizing BM25 and dense kNN methods were conducted using Elasticsearch v8.18. Training and inference for open-source models were carried out on a range of GPUs, including the NVIDIA RTX A6000 Ada, NVIDIA H100, and NVIDIA H200, depending on availability.
For all embedding-based dense retrieval experiments, the questions, generated queries, documents, and summaries were encoded using the OpenAI text-embedding-3-large model, which produces 3072-dimensional vectors.
For fine-tuning experiments with Qwen3-8B, we used the HuggingFace Transformers library Wolf et al. (2020), version 4.51.3, training on the full conversation dataset with a peak learning rate of , a batch size of 8, and a linear learning rate decay schedule.
For inference with open-source models, we utilized vLLM v0.8.5.post1. The sampling hyperparameters used during inference are detailed in Table 11.
| Model Name | top_p | top_k | temperature |
| Gemma3-4B | 0.95 | 64 | 1.0 |
| Qwen3-8B | 0.8 | 20 | 0.7 |
| Qwen3-8B-Think | 0.95 | 20 | 0.6 |
| Qwen3-32B | 0.8 | 20 | 0.7 |
| Qwen3-32B Think | 0.95 | 20 | 0.6 |
| GPT-4.1-mini | 1.0 | - | 1.0 |
| o4-mini | - | - | - |
For query generation in PROBE, we use GPT-4.1-mini as query and filter generator with top_p = 0.5 and top_k = 0.5.
Appendix E Data Construction Process
In this part, we explain in detail how we create the dataset. We start with WildChat-Full dataset which contains around 990K conversations.
E.1 Pre-processing and De-duplication
We begin by de-duplicating the full WildChat dataset using MinHash and Locality-Sensitive Hashing (LSH), following the approach described in Hugging Face (2023). For MinHash, we use 4-grams () and 9 permutations (). For LSH, we set the band size to and the row size to . After de-duplication, approximately 520K conversations remain.
Next, we tokenize all conversations using the LLaMA 3 tokenizer Grattafiori et al. (2024) and discard those exceeding 4,096 tokens. Users are identified based on a combination of hashed IP addresses and HTTP request headers, and each user is assigned a randomized username. Users with fewer than 10 sessions are considered inactive, and all their conversations are removed.
After filtering by conversation length and user activity, around 220K conversations remain. All subsequent processing steps are performed on this filtered dataset.
E.2 LLM-based keywords and summarization extraction
To perform TnT-LLM for topic discovery, we begin by extracting keywords and summaries from raw conversations. Specifically, we prompt GPT-4o to generate both the keyword set and a concise summarization of each conversation. The extracted keywords span a diverse set of semantic types, including persons, technologies, scientific terms, foods, demographic terms, organizations, locations, events, artworks, programming languages, product brands, and financial terms. The complete prompt used for this extraction process is shown in Figure 14.
E.3 TnT-LLM based Topic and Subtopic Discovery and Assignment
E.3.1 Topic Discovery and Assignment
Topic Taxonomy Generation
We largely follow the pipeline of TnT-LLM Wan et al. (2024) to identify topics within the dataset. Rather than randomly sampling from a large corpus, we first obtain the textual embeddings of conversation summaries using the BAAI/bge-en-icl model Li et al. (2024a). We then perform clustering on these embeddings to guide our sampling, ensuring a diverse selection across different semantic regions. This step is added to enhance topic diversity in the sampled subset.
Subsequently, we apply the topic discovery algorithm detailed in Algorithm 1. The initial taxonomy generated is visualized in Figure 15, while the prompt used for topic refinement is shown in Figure 16. For all topic discovery steps, we employ GPT-4o as the underlying language model, using hyperparameters and . To perform efficient KMeans clustering, we utilize the FAISS library Douze et al. (2025). Unlike the original TnT-LLM method, which relies on LLMs for taxonomy refinement, we manually resolve conflicts and enforce mutual exclusivity among the discovered topics.
Topic Label Assignment
Using the generated topics and corresponding taxonomy, we assign a topic ID to each conversation. This assignment process can be formulated as a multi-label classification task. The labeling is performed by GPT-4o using the assignment prompt illustrated in Figure 17. The prompt is carefully designed to mitigate common errors identified through a manual inspection of a small validation set consisting of 400 examples.
E.3.2 Subtopic Discovery and Assignment
Subtopic Taxonomy Generation
For each discovered topic, we further identify its subtopics by running TnT-LLM on all conversations classified under that topic. However, subtopic discovery proves to be more challenging. To address this, we adopt a more sophisticated pipeline and employ a stronger model. The following pipeline is specifically designed to facilitate subtopic discovery within each major topic.
1.
Prompt GPT-4o to check the result of topic assignment and summarize the raw conversation from the perspective of major topic using the prompt shown in Figure 18.
2.
Get the embedding of the summaries that pass checking using text-embedding-3-large.
3.
Run KMeans use faiss with in , find the top 3 best number of centroid using silhouette score Rousseeuw (1987).
4.
For each target number of subtopics ,we execute Algorithm 1 with parameters using topic-specific initial and update prompts as illustrated in Figure 19 and Figure 20. The model used for subtopic discovery is OpenAI-o1, selected for its strong reasoning capabilities. To enforce the desired number of generated subtopics at the start of the iteration, we replace the placeholder “{min_class_number_requirement}” in Figure 19 with instruction “- Generate NO LESS THAN topics.” .
5.
After generating the taxonomy for each , we randomly sample 10% of data instances from the current topic—capped at a maximum of 1000 samples. We then query the o3-mini model, which has strong reasoning ability, using the prompt provided in Figure 21. This yields a set of predicted labels , along with corresponding relevance scores between 0-10, each ranging from 0 to 10. We then compute a quality score for each generated taxonomy using the following equations:
(1) Where and are defined as:
(2) where is the number of samples that labeled as “Undefined”, which is not fit in the taxonomy, and is the number of data sample labeled for taxonomy validation.
(3) We select the best taxonomy generated using .
Subtopic Label Assignment
Finally, we label all data samples using the prompt illustrated in Figure 21, with the o3-mini model. For each topic, we select the best-performing taxonomy and use it to annotate all corresponding samples.
E.4 Topic Label Quality Control
After completing the labeling pipeline, we still observed some false positives upon manual inspection. To address this, we conducted an additional verification step—similar to the initial phase of the subtopic discovery pipeline—by reviewing each data sample alongside its raw conversation, assigned label, and label description, using the o3-mini model and the prompt shown in Figure 22. Following this verification, we removed all samples that lacked a valid label assignment or were assigned the Undefined label at either the topic or subtopic level. This filtering ensured that the final dataset aligned with the discovered taxonomy, ultimately reducing the dataset size to approximately 182k examples.
E.5 Keywords Categorization
After the labeling process, we observed that certain topics—such as “Fanfiction and Crossover” and “Programming” contained a disproportionately large number of data samples. To enable more fine-grained question generation, we further categorized the extracted keywords into four semantic types: programming language, creative artwork, public figure, and book. Conversations that do not contain any keywords from these categories are classified as having no keywords.
E.5.1 LLM Based Aggregation
Assuming that the same word used by the same user conveys a consistent meaning, we first associate each user’s keyword with its corresponding description, extracted at the beginning of the process. We then employ o3-mini to cluster these raw keywords into semantically coherent groups, corresponding to categories including “Programming Language”, “Video Games”, “Tabletop Games”, “Manga/Anime”, “Film”, “TV Show”, “Western Cartoon/Comic”, “Book”, “Musical”, and “Public Figure” , using the prompt illustrated in Figure 23.
E.5.2 Rule-based LLM Result Aggregation
Although o3-mini is prompted to generate the most well-known names for corresponding entities, the model occasionally produces inconsistent outputs, such as “Pokémon” vs. “Pokemon”. These discrepancies are treated as distinct entries in downstream question generation. To address this, we define equivalence between a pair of large language model-generated terms or phrases , where – based on a set of normalization criteria. Terms are considered equivalent across all keyword types except “Public Figure” if they satisfy any of the following conditions after applying string normalization:
1.
and are identical.
2.
and are identical after removing all stopwords in NLTK English stopwords list.
3.
is a prefix of and has more than 2 words.
4.
is a suffix of and has more than 2 words.
5.
is an abbreviation of by concatenating all first letter of .
For keywords of type “Public Figure” only Conditions 1 and 2 are applied due to the higher sensitivity of proper name matching. After normalization, we obtain a dataset with annotated two-level topic hierarchies and keywords spanning the following types: “Programming Language”, “Video Games”, “Tabletop Games”, “Manga/Anime”, “Film”, “TV Show”, “Western Cartoon/Comic”, “Book”, “Musical”, and “Public Figure”.
E.6 Question Proposal
Attributes Combination
We generate questions through a brute-force search over various combinations and quantities of conditions. The full set of considered conditions is shown in Table 1. Specifically, we enumerate all possible attribute combinations containing 0 to 3 conditions and manually select 73 meaningful combinations that can be naturally expressed in language. The selected combinations are listed in Table 8.
Question Proposal Sampling
For each attribute condition and target type combination, we enumerate all possible condition value configurations using MongoDB. For each configuration, we first verify that the number of documents satisfying the condition is at least 50, unless the condition involves the username attribute, in which case the threshold is reduced to 10. This ensures that each generated question is supported by a sufficient number of documents.
Next, we query the database again to check whether the top 3 most frequent target attribute values collectively account for at least 15% of all occurrences. This constraint prevents cases where the target distribution is overly uniform and lacks distinguishing signals.
All condition-target combinations that pass both checks are then stored in a map, where the key is the top-1 target value and the value is a list of corresponding condition-target combinations. Each list is sorted by the normalized entropy of the target distribution to prioritize more informative combinations.
Finally, we sample from this map in a round-robin manner, ensuring that each value is selected no more than twice. This strategy helps generate the most answerable questions while maintaining diversity across different top-1 target outcomes.
E.7 Question Generation
Given a set of condition types, corresponding values, and a target value, we prompt GPT-4.1 to generate natural language questions using the template shown in Figure 24.
Following question generation, we retrieve the top 10 candidate answers for ranking by querying the database. In cases where fewer than 10 valid candidates are available, we supplement them by sampling from the global distribution of values that share the same target type.
Using this procedure, we generated a total of 6,177 questions.
E.8 Question Quality Control
We employ o4-mini for final quality control. Specifically, o4-mini is used to rank target candidates under two settings: (1) without any supporting context, and (2) with supporting context provided in the form of either summaries or raw conversations, using the prompting format shown in Figure 25.For each instance, we compute the instance-wise NDCG@10 score in the no-context setting, denoted as , and define the contextual score as , where and are scores under raw and summarized contexts, respectively.
To assess statistical significance, we calculate a confidence-based threshold to determine whether a contextual improvement is meaningful over random performance. The threshold is defined as:
where is the standard deviation estimated via a Monte Carlo approach, and is the 90%-confidence z-score.We remove any instance that satisfies both of the following conditions:
After filtering, we retain a total of 6,027 valid data samples for downstream evaluation.